OpenStudy (anonymous):

25. Mike's grandmother opened a savings account in Mike's name and deposited some money into the account. The account pays an annual simple interest rate of 3%. After 9 years, the interest earned on the account was $1,350. How much money did Mike's grandmother deposit in the account? A. $5,500 B. $6,000 C. $5,000 D. $4,000

jimthompson5910 (jim_thompson5910):

use the formula I = P*r*t where I = interest earned P = principal (ie amount deposited) r = interest rate t = time in years

jimthompson5910 (jim_thompson5910):

I = P*r*t 1350 = P*0.03*9 solve for P